# Onsager-Machlup Functional and Large Deviation Principle for Stochastic Hamiltonian Systems

### v1

@@ 3. Onsager-Machlup functional for stochastic Hamiltonian systems
In this section, we derive the Onsager-Machlup functional for Hamiltonian systems by calculating the probability ratio of path perturbations within a small neighborhood of a reference path. The main tool applied in this derivation is Girsanov’s theorem. Our result is valid for any finite interval $[0,T]$. However, for simplicity in presentation, we define the interval as $[0,1]$ in the following discussion.
@@
We begin by stating the conditions on the functions $H(q,p)$, $\sigma_{q}(t)$, and $\sigma_{p}(t)$ that will be assumed throughout the proof:
@@
    The Hamiltonian function $H\in C^{3}_{b}(\mathbb{R}^{n}\times\mathbb{R}^{n},\mathbb{R})$, means that $H$ is three times continuously differentiable with bounded third order derivatives. Additionally, the partial derivatives $\frac{\partial H}{\partial q}$ and $\frac{\partial H}{\partial p}$ are globally Lipschitz continuous.
@@
    The diffusion matrices $\sigma_{q}(t)$ and $\sigma_{p}(t)\in C([0,1],\mathbb{R}^{n\times n})$ are positive definite and bounded for any $t\in[0,1]$, and are continuous with respect to $t$. Therefore, the inverses $\sigma_{q}(t)^{-1}$ and $\sigma_{p}(t)^{-1}$ exist, and are continuous and bounded for all $t\in[0,1]$.
@@
(3.1) $\int_{0}^{1}OM(\varphi,\dot{\varphi})\,{\rm d}t=\int_{0}^{1}\left|\dot{\varphi_{q}}(t)-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right|^{2}\,{\rm d}t+\int_{0}^{1}\left|\dot{\varphi_{p}}(t)+\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right|^{2}\,{\rm d}t,$
where $\dot{\varphi}:=\frac{\mathrm{d}\varphi(t)}{\mathrm{d}t}=\left(\frac{\mathrm{d}\varphi_{q}(t)}{\mathrm{d}t},\frac{\mathrm{d}\varphi_{p}(t)}{\mathrm{d}t}\right)$.
@@
Let the reference path be given by $\varphi(t)=(\varphi_{q}(t),\varphi_{p}(t))$, where $\varphi(t)$ is a definite continuous path, and $(\varphi_{q}(t),\varphi_{p}(t))-(q(0),p(0))\in\mathbb{H}^{1}$. We define the perturbed solution, denoted as $(y_{q}(t),y_{p}(t))$, as follows:
(3.2) $\begin{cases}y_{q}(t)=\varphi_{q}(t)+\int_{0}^{t}\sigma_{q}(s)\,{\rm d}W_{q}(s),\\ y_{p}(t)=\varphi_{p}(t)+\int_{0}^{t}\sigma_{p}(s)\,{\rm d}W_{p}(s).\end{cases}$
To simplify the notation in the proof, we introduce the term $W^{\sigma}(t):=\left(W^{\sigma}_{q}(t),W^{\sigma}_{p}(t)\right)$, which represents the stochastic perturbation in the system:
$W^{\sigma}_{q}(t):=\int_{0}^{t}\sigma_{q}(s)\,{\rm d}W_{q}(s),\quad W^{\sigma}_{p}(t):=\int_{0}^{t}\sigma_{p}(s)\,{\rm d}W_{p}(s).$
We define $\tilde{W}_{q}(t)$ and $\tilde{W}_{p}(t)$ as follows. It can be shown that under the new probability measures ~$\tilde{\mathbb{P}}_{q}$ and ~$\tilde{\mathbb{P}}_{p}$, $\tilde{W}_{q}(t)$ and $\tilde{W}_{p}(t)$ are standard Brownian motions.
(3.3) $\tilde{W}_{q}(t)$ $=W_{q}(t)-\int_{0}^{t}\sigma_{q}^{-1}(s)\left(\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})-\dot{\varphi}_{q}(s)\right)\,{\rm d}s,$
$\tilde{W}_{p}(t)$ $=W_{p}(t)-\int_{0}^{t}\sigma_{p}^{-1}(s)\left(-\frac{\partial H}{\partial y_{q}}(y_{q},y_{p})-\dot{\varphi}_{p}(s)\right)\,{\rm d}s.$
@@
(3.4) $\begin{cases}\,{\rm d}y_{q}(t)=\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})dt+\sigma_{q}(t)\,{\rm d}\tilde{W}_{q}(t),\\ \,{\rm d}y_{p}(t)=-\frac{\partial H}{\partial y_{q}}(y_{q},y_{p})dt+\sigma_{p}(t)\,{\rm d}\tilde{W}_{p}(t).\end{cases}$
@@
To apply Girsanov’s Theorem and achieve the transformation between the two measures, we define the Radon-Nikodym derivative $\mathcal{R}:=\frac{d\tilde{\mathbb{P}}}{d\mathbb{P}}=\frac{{\rm d}\tilde{\mathbb{P}}_{q}}{{\rm d}\mathbb{P}_{q}}\cdot\frac{{\rm d}\tilde{\mathbb{P}}_{p}}{{\rm d}\mathbb{P}_{p}}$, which represents the change of measure from $\mathbb{P}$ to $\tilde{\mathbb{P}}$. This derivative is given by an exponential martingale associated with the drift terms, which describes the behavior of the Brownian motion under the new measure after the removal of the drift. For the position variable $q$, the Radon-Nikodym derivative is:
$\frac{d\tilde{\mathbb{P}_{q}}}{d\mathbb{P}_{q}}=\exp$ $\left(\int_{0}^{1}\left\langle\sigma_{q}^{-1}(s)\left(\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})-\dot{\varphi}_{q}(s)\right),\,{\rm d}W_{q}(s)\right\rangle\right.$
$\left.-\frac{1}{2}\int_{0}^{1}\left|\sigma_{q}^{-1}(s)\left(\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})-\dot{\varphi}_{q}(s)\right)\right|^{2}\,{\rm d}s\right),$
and similarly for the momentum variable $p$:
$\frac{d\tilde{\mathbb{P}}_{p}}{d\mathbb{P}_{p}}=\exp$ $\left(\int_{0}^{1}\left\langle\sigma_{p}^{-1}(s)\left(-\frac{\partial H}{\partial y_{q}}(y_{q},y_{p})-\dot{\varphi}_{p}(s)\right),\,{\rm d}W_{p}(s)\right\rangle\right.$
$\left.-\frac{1}{2}\int_{0}^{1}\left|\sigma_{p}^{-1}(s)\left(\frac{\partial H}{\partial y_{q}}(y_{q},y_{p})+\dot{\varphi}_{p}(s)\right)\right|^{2}\,{\rm d}s\right).$
So,
$\mathcal{R}=\exp$ $\left(\int_{0}^{1}\left\langle\sigma_{q}^{-1}(s)\left(\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})-\dot{\varphi}_{q}(s)\right),\,{\rm d}W_{q}(s)\right\rangle\right.$
$-\int_{0}^{1}\left\langle\sigma_{p}^{-1}(s)\left(\frac{\partial H}{\partial y_{q}}(y_{q},y_{p})-\dot{\varphi}_{p}(s)\right),\,{\rm d}W_{p}(s)\right\rangle$
$-\frac{1}{2}\int_{0}^{1}\left|\sigma_{q}^{-1}(s)\left(\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})-\dot{\varphi}_{q}(s)\right)\right|^{2}\,{\rm d}s$
$\left.-\frac{1}{2}\int_{0}^{1}\left|\sigma_{p}^{-1}(s)\left(\frac{\partial H}{\partial y_{q}}(y_{q},y_{p})+\dot{\varphi}_{p}(s)\right)\right|^{2}\,{\rm d}s\right).$
We now aim to compute the transition probability of the system’s path remaining close to the reference path $\varphi(t)$. Using Girsanov’s Theorem, this probability can be expressed as:
(3.5) $\quad\frac{\mathbb{P}\left(\left\|(q,p)-(\varphi_{q},\varphi_{p})\right\|\leq\epsilon\right)}{\mathbb{P}\left(\left\|W^{\sigma}\right\|\leq\epsilon\right)}=\frac{\tilde{\mathbb{P}}\left(\left\|(Y_{q},Y_{p})-(\varphi_{q},\varphi_{p})\right\|\leq\epsilon\right)}{\mathbb{P}\left(\left\|W^{\sigma}\right\|\leq\epsilon\right)}$
=~$=\frac{\tilde{\mathbb{P}}\left(\left\|W^{\sigma}\right\|\leq\epsilon\right)}{\mathbb{P}\left(\left\|W^{\sigma}\right\|\leq\epsilon\right)}=\frac{\mathbb{E}\left(\mathcal{R}\mathbb{I}_{\left\|W^{\sigma}\right\|\leq\epsilon}\right)}{\mathbb{P}\left(\left\|W^{\sigma}\right\|\leq\epsilon\right)}=\mathbb{E}\left(\mathcal{R}\big{|}\left\|W^{\sigma}\right\|\leq\epsilon\right)$
$=\exp\left\{-\frac{1}{2}\left(\int_{0}^{1}\left|\sigma_{q}^{-1}(t)\left(\dot{\varphi}_{q}-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right)\right|^{2}\,{\rm d}t\right.\right.$
$\qquad\qquad\quad\left.\left.+\int_{0}^{1}\left|\sigma_{p}^{-1}(t)\left(\dot{\varphi}_{p}+\frac{\partial H}{\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\right)\right|^{2}\,{\rm d}t\right)\right\}$
×$\quad\times\mathbb{E}\left(\exp\left\{\sum_{i=1}^{6}B_{i}\right\}\bigg{|}\left\|W^{\sigma}\right\|\leq\epsilon\right),$
@@
$B_{1}$ $=\int_{0}^{1}\left\langle\sigma_{q}^{-1}(s)\frac{\partial H}{\partial y_{p}}(y_{q},y_{p}),\,{\rm d}W_{q}(s)\right\rangle-\int_{0}^{1}\left\langle\sigma_{p}^{-1}(s)\frac{\partial H}{\partial y_{q}}(y_{q},y_{p}),\,{\rm d}W_{p}(s)\right\rangle,$
$B_{2}$ $=-\int_{0}^{1}\left\langle\sigma_{q}^{-1}(s)\dot{\varphi}_{q}(s),\,{\rm d}W_{q}(s)\right\rangle-\int_{0}^{1}\left\langle\sigma_{p}^{-1}(s)\dot{\varphi}_{p}(s),\,{\rm d}W_{p}(s)\right\rangle,$
$B_{3}$ $=\frac{1}{2}\int_{0}^{1}\left|\sigma_{q}^{-1}(s)\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right|^{2}\,{\rm d}s-\frac{1}{2}\int_{0}^{1}\left|\sigma_{q}^{-1}(s)\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})\right|^{2}\,{\rm d}s,$
$B_{4}$ $=\frac{1}{2}\int_{0}^{1}\left|\sigma_{p}^{-1}(s)\frac{\partial H}{\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\right|^{2}\,{\rm d}s-\frac{1}{2}\int_{0}^{1}\left|\sigma_{p}^{-1}(s)\frac{\partial H}{\partial y_{q}}(y_{q},y_{p})\right|^{2}\,{\rm d}s,$
$B_{5}$ $=\int_{0}^{1}\left\langle\sigma_{q}^{-2}(s)\left(\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right),\dot{\varphi}_{q}(s)\right\rangle\,{\rm d}s,$
$B_{6}$ $=-\int_{0}^{1}\left\langle\sigma_{p}^{-2}(s)\left(\frac{\partial H}{\partial y_{q}}(y_{q},y_{p})-\frac{\partial H}{\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\right),\dot{\varphi}_{p}(s)\right\rangle\,{\rm d}s.$
For the second term $B_{2}$, we have
$B_{2}$ $=-\int_{0}^{1}\left\langle\sigma_{q}^{-1}(s)\dot{\varphi}_{q}(s),\,{\rm d}W_{q}(s)\right\rangle-\int_{0}^{1}\left\langle\sigma_{p}^{-1}(s)\dot{\varphi}_{p}(s),\,{\rm d}W_{p}(s)\right\rangle$
$=-\sum_{i=1}^{n}\left(\int_{0}^{1}\sigma_{q,i}^{-1}(s)\dot{\varphi}_{q,i}(s)\,\,{\rm d}W_{q,i}(s)+\int_{0}^{1}\sigma_{p,i}^{-1}(s)\dot{\varphi}_{p,i}(s)\,\,{\rm d}W_{p,i}(s)\right).$
It is straightforward to demonstrate that $\sigma_{q,i}^{-1}(s)\dot{\varphi}_{q,i}(s)\in L^{2}$ and $\sigma_{p,i}^{-1}(s)\dot{\varphi}_{p,i}(s)\in L^{2}$ for all $0<i<n$. Applying Lemma 2.8 and Lemma 2.13, we subsequently obtain
(3.6) $\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{2}\right\}\big{|}\left\|W^{\sigma}\right\|<\epsilon\right)=1$
@@
For the third term $B_{3}$,
$B_{3}$ $=\frac{1}{2}\int_{0}^{1}\left|\sigma_{q}^{-1}(s)\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right|^{2}\,{\rm d}s-\frac{1}{2}\int_{0}^{1}\left|\sigma_{q}^{-1}(s)\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})\right|^{2}\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}{\sigma_{q}^{-2}(s)\left|\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})-\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})\right|^{2}}$
$\quad+{2\sigma_{q}^{-2}(s)\left|\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})-\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})\right|\left|\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})\right|}\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}{\sigma_{q}^{-2}(s)\left|\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})-\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})\right|^{2}\,{\rm d}s}$
$\quad+\int_{0}^{1}{\sigma_{q}^{-2}(s)\left|\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})-\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})\right|\left|\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})\right|}\,{\rm d}s.$
In Condition $C1$, since $\frac{\partial H}{\partial p}$ is Lipschitz continuous, we have the following estimate:
(3.7) $\quad\left|\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right|$
$=\left|\frac{\partial H}{\partial\left(\varphi_{p}+W^{\sigma}_{p}\right)}((\varphi_{q}+W^{\sigma}_{q}),(\varphi_{p}+W^{\sigma}_{p}))-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right|\leq L\left\|W^{\sigma}\right\|.$

### v2

@@ 3 Onsager-Machlup functional for stochastic Hamiltonian systems
In this section, we derive the Onsager-Machlup functional for Hamiltonian systems by calculating the probability ratio of path perturbations within a small neighborhood of a reference path. The main tools used in the derivation are Girsanov’s theorem, the conditional expectation comparison principle, and the results from Sections 2.4 and 2.5. Our results hold for any finite interval $[0,T]$. However, for simplicity, we define the interval as $[0,1]$ in the following discussion. It is worth noting that, to the best of our knowledge, this is the first Onsager-Machlup functional for multiplicative noise in Euclidean space.
@@
We begin by stating the conditions on the functions $H(\theta,I)$ and $\sigma(t,\theta,I)$ that will be assumed throughout the proof:
@@
    The Hamiltonian function $H\in C^{3}_{b}(\mathbb{T}^{n}\times\mathbb{R}^{n},\mathbb{R})$ is three times continuously differentiable with all third-order derivatives bounded. Furthermore, its gradient $\nabla H$ is globally Lipschitz continuous with a Lipschitz constant $L$.
@@
    The diffusion matrix $\sigma\in C^{2}_{b}([0,1]\times\mathbb{T}^{n}\times\mathbb{R}^{n},\mathbb{R}^{n\times n})$ is symmetric, positive definite, and bounded for all $t\in[0,1]$. Moreover, it is Lipschitz continuous with respect to the second and third variables, with Lipschitz constant $L$. In addition, $\sigma$ satisfies a uniform ellipticity condition: there exist constants $0<\lambda\leq\Lambda$ such that
    $\lambda\leq\|\sigma\|_{HS}\leq\Lambda,\quad\forall t\in[0,1],$
    where $\|\cdot\|_{HS}$ denotes the Hilbert–Schmidt norm. Consequently, the inverse $\sigma^{-1}$ exists and is continuous, bounded, and symmetric for all $t\in[0,1]$.
@@
$OM(\varphi,\dot{\varphi})$ $=\int_{0}^{1}{\left|\sigma^{-1}(t,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi_{s}}\right)\right|^{2}}\,{\rm d}s-\int_{0}^{1}\mathrm{Tr}\left(\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right)\,{\rm d}s,$ (8)
where $\dot{\varphi}(t):=\left(\dot{\varphi_{\theta}}(t),\dot{\varphi_{I}}(t)\right)$, and $|\cdot|$ denotes the Euclidean norm in $\mathbb{R}^{n}$.
@@
Let the reference path be given by $\varphi(t)=(\varphi_{\theta}(t),\varphi_{I}(t))$, where $\varphi(t)$ is a definite continuous path, and $\varphi(t)-X(0)\in\mathbb{H}^{1}$. We define the stochastic perturbed solution, denoted as $y(t):=(y_{\theta}(t),y_{I}(t))$, as follows:
$y(t)=\varphi(t)+\int_{0}^{t}\sigma(s,y_{s})\,{\rm d}W_{s}.$ (9)
To simplify the notation in the proof, we introduce the term $W^{\sigma}(t)$, which represents the stochastic perturbation in the system:
$W^{\sigma}(t):=\int_{0}^{t}\sigma(s,y_{s})\,{\rm d}W_{s}.$
We define $\tilde{W}(t)$ as follows. Since $\sigma$ is a predictable process, it can be shown, using Girsanov’s Theorem [24], that under the new probability measure $\tilde{\mathbb{P}}$, the process $\tilde{W}(t)$ is a standard $2n$-dimensional Brownian motion.
$\tilde{W}(t)$ $=W(t)-\int_{0}^{t}\sigma^{-1}(s,y_{s})\left(J\nabla H(y_{s})-\dot{\varphi}_{s}\right)\,{\rm d}s.$ (10)
@@
$\mathrm{d}y(t)=J\nabla H(y_{t})\,\mathrm{d}t+\sigma(t,y_{t})\,\mathrm{d}\tilde{W}_{t},$ (11)
@@
To apply Girsanov’s Theorem and achieve the transformation between the two measures, we define the Radon-Nikodym derivative $\mathcal{R}:=\frac{d\tilde{\mathbb{P}}}{d\mathbb{P}}$, which represents the change of measure from $\mathbb{P}$ to $\tilde{\mathbb{P}}$. This derivative is given by an exponential martingale associated with the drift terms, which describes the behavior of the Brownian motion under the new measure after the removal of the drift. The Radon-Nikodym derivative is
$\frac{d\tilde{\mathbb{P}}}{d\mathbb{P}}=\exp$ $\left(\int_{0}^{1}\left\langle\sigma^{-1}(s,y_{s})\left(J\nabla H(y_{s})-\dot{\varphi}_{s}\right),\,{\rm d}W_{s}\right\rangle-\frac{1}{2}\int_{0}^{1}\left|\sigma^{-1}(s,y_{s})\left(J\nabla H(y_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s\right),$
We now aim to compute the transition probability of the system’s path remaining close to the reference path $\varphi(t)$. Using Girsanov’s Theorem, this probability can be expressed as
$\quad\frac{\mathbb{P}\left(\left\|X-\varphi\right\|\leq\epsilon\right)}{\mathbb{P}\left(\left\|W^{\sigma}\right\|\leq\epsilon\right)}=\frac{\tilde{\mathbb{P}}\left(\left\|y-\varphi\right\|\leq\epsilon\right)}{\mathbb{P}\left(\left\|W^{\sigma}\right\|\leq\epsilon\right)}=\frac{\tilde{\mathbb{P}}\left(\left\|W^{\sigma}\right\|\leq\epsilon\right)}{\mathbb{P}\left(\left\|W^{\sigma}\right\|\leq\epsilon\right)}$ (12)
$=\frac{\mathbb{E}\left(\mathcal{R}\mathbb{I}_{\left\|W^{\sigma}\right\|\leq\epsilon}\right)}{\mathbb{P}\left(\left\|W^{\sigma}\right\|\leq\epsilon\right)}=\mathbb{E}\left(\mathcal{R}\big{|}\left\|W^{\sigma}\right\|\leq\epsilon\right)$
$=\exp\left\{-\frac{1}{2}\left(\int_{0}^{1}\left|\sigma^{-1}(t,\varphi_{s})\left(\dot{\varphi_{s}}-J\nabla H(\varphi_{s})\right)\right|^{2}\,{\rm d}s\right)\right\}\times\mathbb{E}\left(\exp\left\{\sum_{i=1}^{4}B_{i}\right\}\bigg{|}\left\|W^{\sigma}\right\|\leq\epsilon\right),$
@@
$B_{1}$ $=\int_{0}^{1}\left\langle\sigma^{-1}(s,y_{s})J\nabla H(y_{s}),\,{\rm d}W_{s}\right\rangle,$
$B_{2}$ $=-\int_{0}^{1}\left\langle\sigma_{\theta}^{-1}(s,y_{s})\dot{\varphi}_{s},\,{\rm d}W_{s}\right\rangle,$
$B_{3}$ $=\frac{1}{2}\int_{0}^{1}\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(y_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s-\frac{1}{2}\int_{0}^{1}\left|\sigma^{-1}(s,y_{s})\left(J\nabla H(y_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s,$
$B_{4}$ $=\frac{1}{2}\int_{0}^{1}\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s-\frac{1}{2}\int_{0}^{1}\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(y_{s})-\dot{\varphi}_{s}\right)\right|^{2}\,{\rm d}s.$
For the first term $B_{1}$, expand $F(t,y):=\sigma^{-1}(t,y_{t})J\nabla H(y_{t})$ in a Taylor series around $x=\varphi_{t}$ with respect to $W^{\sigma}_{t}:=W^{\sigma}(t)$. The expansion is given by
$F(t,y_{t})=F(t,\varphi)+D_{y}F(t,\varphi)\cdot\left(y-\varphi\right)+O(\|y-\varphi\|^{2}).$
where $D_{y}F\in\mathbb{R}^{2n\times 2n}$ is the Jacobian matrix of $F$ with respect to $y$, and $O(\|y-\varphi\|^{2})$ denotes the higher-order remainder term. The specific expansion for $B_{1}$ can be written as
$B_{1}$ $=\int_{0}^{1}\left\langle\sigma^{-1}(s,y_{s})J\nabla H(y_{s}),\,{\rm d}W_{s}\right\rangle$
$=\int_{0}^{1}\left\langle{\left(\sigma^{-1}(s,\varphi_{s})J\nabla H(\varphi_{s})+D_{y}\left(\sigma^{-1}(s,\varphi_{s})J\nabla H(\varphi_{s})\right)W^{\sigma}_{s}+R(s)\right)},\,{\rm d}W_{s}\right\rangle$
$=B_{11}+B_{12}+B_{13},$
where:
$\mathbb{E}\left|R(t)\right|^{2}=\mathcal{O}\left(\mathbb{E}\|W^{\sigma}_{t}\|^{4}\right)=\mathcal{O}(\epsilon^{4}),\quad\text{for }\|A\|\leq\epsilon.$
For the first term $B_{2}$, we proceed with a similar Taylor expansion for $B_{1}$
$B_{2}$ $=-\int_{0}^{1}\left\langle\sigma^{-1}(s,y_{s})\dot{\varphi}_{s},\,{\rm d}W_{s}\right\rangle$
$=-\int_{0}^{1}\left\langle{\left(\sigma^{-1}(s,\varphi_{s})\dot{\varphi}_{s}+D_{y}\left(\sigma^{-1}(s,\varphi_{s})\dot{\varphi}_{s}\right)W^{\sigma}_{s}+R(s)\right)},\,{\rm d}W_{s}\right\rangle$
$=B_{21}+B_{22}+B_{23}.$
Since $B_{11}$ and $B_{21}$, $B_{12}$ and $B_{22}$, as well as $B_{13}$ and $B_{23}$ exhibit similar structural forms, we analyze each pair jointly in the subsequent discussion. Based on conditions $(C_{1})$, $(C_{2})$, and $\varphi(t)-X(0)\in\mathbb{H}^{1}$, we can demonstrate that the function
$\sigma^{-1}(t,\varphi_{t})J\nabla H(\varphi_{t})~\text{and}~\sigma^{-1}(t,\varphi_{t})\dot{\varphi}_{t}\in L^{2}([0,1],\mathbb{R}^{2n}).$
Therefore, according to Lemma 2.20, we have
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{11}\right\}\big{|}\|W^{\sigma}\|\leq\epsilon\right)=1$ (13)
for all $c\in\mathbb{R}$. And
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{21}\right\}\big{|}\|W^{\sigma}\|\leq\epsilon\right)=1$ (14)
@@
Secondly, by applying the matrix version of the chain rule for differentiation, we obtain
$B_{12}$ $=\int_{0}^{1}\left\langle D_{y}\left(\sigma^{-1}(s,\varphi_{s})J\nabla H(\varphi_{s})\right)W^{\sigma}_{s},\,{\rm d}W_{s}\right\rangle$
$=\int_{0}^{1}\left\langle D_{y}\left(\sigma^{-1}(s,\varphi_{s})\right)J\nabla H(\varphi_{s})W^{\sigma}_{s}+\sigma^{-1}(s,\varphi_{s})JD_{y}\nabla H(\varphi_{s})W^{\sigma}_{s},\,{\rm d}W_{s}\right\rangle$
$=\int_{0}^{1}\left\langle-\sigma^{-1}(s,\varphi_{s})\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})J\nabla H(\varphi_{s})W^{\sigma}_{s}\right$.
$\quad\left.+\sigma^{-1}(s,\varphi_{s})JD_{y}\nabla H(\varphi_{s})W^{\sigma}_{s},\,{\rm d}W_{s}\right\rangle,$
where $D_{y}\sigma(t,\varphi_{t})$ refers to the derivative tensor of the matrix-valued function $\sigma(t,\varphi_{t})$ with respect to $y_{t}\in\mathbb{R}^{n}$. Specifically, $D_{y}\sigma(t,\varphi_{t})$ is a third-order tensor of dimension $n\times n\times n$, whose components are given by
$(D_{y}\sigma(t,\varphi_{t}))_{ijk}:=\frac{\partial\sigma_{ij}(t,y_{t})}{\partial y_{k}}\bigg{|}_{(t,\varphi_{t})}\quad\text{for }i,j,k=1,\dots,n,$
where $\sigma_{ij}(t,y_{t})$ denotes the $(i,j)$-th element of the matrix $\sigma(t,y_{t})$, and $\frac{\partial\sigma_{ij}(t,y_{t})}{\partial y_{k}}$ is the partial derivative of $\sigma_{ij}(t,y_{t})$ with respect to the $k$-th component of $y_{t}$ (i.e., the different coordinates of $y_{k}$). And
$D_{y}\nabla H(\varphi_{t})=\begin{pmatrix}\frac{\partial^{2}H}{\partial\varphi_{\theta}^{2}}&\frac{\partial^{2}H}{\partial\varphi_{\theta}\partial\varphi_{I}}\\ \frac{\partial^{2}H}{\partial\varphi_{I}\partial\varphi_{\theta}}&\frac{\partial^{2}H}{\partial\varphi_{I}^{2}}\end{pmatrix}$
represents the Hessian matrix of $H$. And
$B_{22}$ $=-\int_{0}^{1}\left\langle D_{y}\left(\sigma^{-1}(s,\varphi_{s})\dot{\varphi}_{s}\right)W^{\sigma}_{s},\,{\rm d}W_{s}\right\rangle$
$=\int_{0}^{1}\left\langle\sigma^{-1}(s,\varphi_{s})\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\dot{\varphi}_{s}W^{\sigma}_{s},\,{\rm d}W_{s}\right\rangle.$
To simplify the notation, define $F(t,\varphi_{t})\in\mathbb{R}^{n\times n}$ as
$F(t,\varphi_{t})=\sigma^{-1}(t,\varphi_{t})\left[JD_{y}\nabla H(\varphi_{t})-\left(D_{y}\sigma(t,\varphi_{t})\right)\sigma^{-1}(t,\varphi_{t})\left(J\nabla H(\varphi_{t})-\dot{\varphi}_{t}\right)\right].$
Hence, by substituting the expression for $W^{\sigma}(t)$, we obtain
$B_{12}+B_{22}$ $=\int_{0}^{1}\left\langle{F(s,\varphi_{s})W^{\sigma}_{s}},\,{\rm d}W_{s}\right\rangle$
$=\int_{0}^{1}\left\langle{F(s,\varphi_{s})\int_{0}^{s}{\sigma\left(t,y_{t}\right)}\,{\rm d}W_{t}},\,{\rm d}W_{s}\right\rangle$
$=\int_{0}^{1}\left\langle{F(s,\varphi_{s})\int_{0}^{s}{\sigma\left(t,\varphi_{t}\right)}\,{\rm d}W_{t}},\,{\rm d}W_{s}\right\rangle$
$\quad+\int_{0}^{1}\left\langle{F(s,\varphi_{s})\int_{0}^{s}\left({\sigma\left(t,y_{t}\right)-\sigma\left(t,\varphi_{t}\right)}\right)\,{\rm d}W_{t}},\,{\rm d}W_{s}\right\rangle$
$=B_{\cdot 21}+B_{\cdot 22}.$
We will examine the term $B_{\cdot 21}$, which can be expressed as a double stochastic integral with respect to $W$:
$B_{\cdot 21}$ $=\int_{0}^{1}\int_{0}^{s}\left\langle{F(s,\varphi_{s})\sigma\left(t,\varphi_{t}\right)}\,{\rm d}W_{t},\,{\rm d}W_{s}\right\rangle.$
Define the function $\bar{F}(s,t)$ as follows:
$\bar{F}(s,t):=F(s,\varphi_{s})\sigma(t,\varphi_{t}).$
Additionally, define the symmetrized matrix of $\bar{F}$, denoted by $\tilde{F}$, as
$\tilde{F}:=\frac{1}{2}(\bar{F}+\bar{F}^{*}).$
Thus, we have
$B_{\cdot 21}=\int_{0}^{1}\int_{0}^{1}\left\langle\tilde{F}\,{\rm d}W_{t},\,{\rm d}W_{s}\right\rangle.$
According to condition $(C2)$, the integral operator K(F~)K($\tilde${F}) associated with the symmetrized kernel $\tilde{F}$ is nuclear, and its trace can be computed as follows: